I need assistance on completing the square in order graph the conic section.
-9x^2+4y^2+36x+24y-36=0
Do you start off as -9x^2+4y^2+36x+24y=-36?

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(9144)

first put all the x's together and all the y's together
then move the constant -36 to the other side
then do the complete the square calculations first for x and then for y

Answer by Alan3354(31513)

-9x^2+4y^2+36x+24y-36=0
Do you start off as -9x^2+4y^2+36x+24y=-36
----------------------
That's a good step 1.
Group the x and y terms:
-9x^2+36x + 4y^2+24y = -36
-(9x^2-36x) + 4y^2+24y = -36
Add coeffs to complete the squares:
-(9x^2-36x+36) + (4y^2+24y+36) = -36 - 36 + 36
-(3x-6)^2 + (2y+6)^2 = -36
(2y+6)^2 - (3x-6)^2 = 36